OpenStudy (anonymous):

How would I verify the equation: cot^4(x)+2 cot^2(x)+1=csc^4(x)

OpenStudy (anonymous):

Would I factor it or attempt to?

ganeshie8 (ganeshie8):

hint : a^2+2ab+b^2 = (a+b)^2

ganeshie8 (ganeshie8):

hint2 : 1 = 1^2

OpenStudy (anonymous):

Can I us x's in replacement of cot and replace cot later?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

So the factored eq would be:\[(\cot^2\theta +1)^2=\csc^4\theta ?\]

ganeshie8 (ganeshie8):

looks good ^

OpenStudy (anonymous):

and then solve from there?

ganeshie8 (ganeshie8):

you're done ! there is nothing to solve actually :)

OpenStudy (anonymous):

wow...

ganeshie8 (ganeshie8):

just use the identity \(\large 1 + \cot ^2 \theta = \csc ^2 \theta\)

OpenStudy (anonymous):

Oh then use trig identities to say that^^^

ganeshie8 (ganeshie8):

\(\large \cot^4(x)+2 \cot^2(x)+1\) say cot^2 = x \(\large x^2 + 2x+1\) \(\large (x+1)(x+1)\) \(\large (x+1)^2\)

ganeshie8 (ganeshie8):

plugin x = cot^2 and then use the trig identity

OpenStudy (anonymous):

Thank you so much. This community is as good as a school teacher. going into college this year and am trying to remember most of the trig stuff.

ganeshie8 (ganeshie8):

\(\large ( \cot^2 \theta + 1)^2\) \(\large ( \csc^2\theta )^2\) \(\large \csc^4 \theta\) which is same as right hand side. so we're done !

ganeshie8 (ganeshie8):

np, you're welcome :) you must be excited about the college life ! good luck !!

OpenStudy (anonymous):

ty